the electric potential decreases uniformly from 120V to 80V as one moves on x axis from x=-1 to x=+1.then electric field at origin is 
options
must be equal to 20V/cm 
may be equal to to 20V/vm
may be greater than 20V/cm
maybe lessthan 20V/cm

Grade:9

3 Answers

Arun
6 years ago
E = -dV/dx 
Since V decreases uniformly, E = -ΔV/Δx 
= -(80 - 120)/{1- (-1)} 
= - ( -40)/2 
= 20 V/Cm in the positive direction of x axis.
Hence option 2 and 3 are correct.
Regards
Arun (askIITians forum expert)
 
Brijesh Rohit
5 years ago
E=-dV/dr gives the component of the electric field in the direction of displacement  dr.
Thus 20V/cm is the component of electric field  in the direction of x axis which means electric field can be equal to 20V/cm or greater than that but it's component will be equal to 20V/cm in the direction of x axis. 
Kushagra Madhukar
3 years ago
As we know,
E = – dV/dx =-(V2-V1)/(x2-x1)  (since V varies uniformly)
E = – (80 – 120)/(1-[-1])
    = 40/2 = 20 V/cm
 
Hope it helps
Regards,
Kushagra